From GlueXWiki
Jump to: navigation, search

October 11, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Corrosion tests (Vlad)
  3. Engineering Ansys analyses of the cooling (Bill)
  4. Electronics (Chris, Nick)
  5. Third package tests FDC E-log (Beni)
  6. Other


Participants: Eugene, Bill, Dave, Chris, Nick, Vlad, Simon, Beni, and Lubomir.


- Dave: The the first package was assembled completely on Friday. Due to delays with the gas bottles (again GTS company!), we started flushing with Nitrogen. Will have the bottles by the end of the next week hopefully. Anatoly is working on the cathodes for package #4. All but one are ready. Wire frame #29 will be deadened on Monday. Need Casey and Mike for several hours to string wire frame #30. These frames are needed for package #4 as well. PO for the foils was issued yesterday, expect them Nov 15, then Anatoly will start testing them. We need the foils for package #2 sometime in December.

- Vlad started applying HV on the new cells: could not reach more than ~1800V on two of the new cells. We suspect that the limitation comes from the gas, Nitrogen. We will try the same on a cell that worked before.

Cathode corrosion

- Vlad showed new results of the tests in the vacuum chamber at high temperature (~100degC), high humidity (~100%) in air. One of the samples showed >1MOhm after 4 days, while the other two are OK after 11 days. We have three samples, EPDM, Viton, and Viton+Apiezon, but the experiment was blinded, so we don't know which is which.


- Bill made ANSYS analyses of the cooling (document attached). Conclusion: the parallel supply as was designed before is good enough. The flows through different loops differ at most few percents. The temperature difference along one loop is less tha 2degC. Bill will go ahead with purchasing the Temperature Control Unit. He submitted also a PR for the fiducial hardware.


- Chris: PR for the ground clips went through, it will take a month from now to get the clips. PR for the daughter card population not awarded yet, will have all the parts by tomorrow, the turnaround time is a week.

Third package tests

- Beni and Lubomir looked at the two "missing wires" with a scope. One turned out to be a channel of the pre-amp not working, the other one was a result of the pre-amp card not completely connected, due to a slight angle of the cooling tube, in turn, caused by a cathode card glued slightly off radially. Now we see signals from all the wires.

- We also looked with an iron source using scope, analog card+fADC and discriminator card+F1TDC, in the last two cases using some strips for triggering. At 2150/-500V very small percentage (2-3%) after-pulsing was observed. We conclude the hit multiplicity that Beni reported (Elog #90) is caused by the cosmic clustering itself. To check this Lubomir did Garfield simulations (Elog #92) that explain the hit multiplicity, but still they don't show the long tail in distribution of the timing between the first hit and all the successive ones (Elog #92).

- Beni now reports ~100% wire efficiency (Elog #91) with a plateau starting at 2050V measured up to 2150V, but can be extended up to 2300V if we have to (Lubomir). Eugene: one has to measure also the noise and put it on the plateau plot.

- Simon is looking into the hddm data from the cosmic tests. Having only wire information from the six FDC cells he will try to reconstruct straight tracks and look at the residuals.

- Temperature at 126 is still very high (80s F), Eugene is contacting the responsible people and the answer so far: within the present HVAC system in the ELL there's nothing that can be done. Bill agreed to look for different solutions like installing additional cooling/ventilation/openings in the wall.